OpenStudy (anonymous):

TRIGONOMETRY through what angle in degrees must the pendulum with 40cm. long be rotated to travel an arclength of 5cm.? TRIGONOMETRY through what angle in degrees must the pendulum with 40cm. long be rotated to travel an arclength of 5cm.? @Mathematics

OpenStudy (anonymous):

arclength=(2*pi*r*angle/360) ==>angle=(360*arclength/(2*pi*r))=(360*5/(2*pi*40))=7.16 degrees

OpenStudy (anonymous):

5/40 radian=.125 radian